How do I find the radius of a this cylinder?
The height (h) is twice the radius (r) The total area is 150π
I set up the equation 2πr(2r)+2πr^2 = 150π
now what?

Answer by richwmiller(17219)

what is the formula for the area of a cylinder?
Does total area include the top and bottom surface areas?
height =2r radius =r area =150
2 pi r^2 h is the lateral surface area
2 pi r (r+h) all surface areas including top and bottom
2 pi r^2*2r =150 pi
4 pi r^3=150
r^3=150/4
r^3=75
r=3.35
or
2 pi r (r+h)=150 pi
2 pi r *(3r)=150 pi
6 pi r^3=150 pi
r^3=150/6
r^3=25
r=2.92